# Generative AI for Data Augmentation in Wireless Networks: Analysis, Applications, and Case Study

### v1

@@ Abstract
Data augmentation is a powerful technique to mitigate data scarcity. However, owing to fundamental differences in wireless data structures, traditional data augmentation techniques may not be suitable for wireless data. Fortunately, Generative Artificial Intelligence (GenAI) can be an effective alternative to wireless data augmentation due to its excellent data generation capability. This article systemically explores the potential and effectiveness of GenAI-driven data augmentation in wireless networks. We first briefly review data augmentation techniques, discuss their limitations in wireless networks, and introduce generative data augmentation, including reviewing GenAI models and their applications in data augmentation. We then explore the application prospects of GenAI-driven data augmentation in wireless networks from the physical, network, and application layers, which provides a GenAI-driven data augmentation architecture for each application. Subsequently, we propose a general generative diffusion model-based data augmentation framework for Wi-Fi gesture recognition, which uses transformer-based diffusion models to generate high-quality channel state information data. Furthermore, we develop residual neural network models for Wi-Fi gesture recognition to evaluate the role of augmented data and conduct a case study based on a real dataset. Simulation results demonstrate the effectiveness of the proposed framework. Finally, we discuss research directions for generative data augmentation.
@@
Wireless networks, data augmentation, GenAI, Wi-Fi gesture recognition, diffusion models.
@@ I Introduction
As an indispensable and fundamental technology, wireless networks enable users to access the Internet, connect devices, and communicate wirelessly. The growing complexity and diversity of wireless networks have facilitated the development of Deep Learning (DL)-based wireless communications and networking[1]. Since wireless networks generate large amounts of data from user activities, channel states, and network conditions, DL can efficiently distill high-level features and information from these data, which have complex structures and inner correlations[1], and enable intelligent applications such as Digital Twins (DTs), autonomous driving, and metaverse. Nevertheless, the availability of vast amounts of high-quality wireless data is a major determinant of the effectiveness of DL models[2]. Specifically,
  • Dynamic wireless channels: Due to multipath fading, shadowing, and interference, the environment in wireless channels is highly variable[3], which makes it difficult to collect high-quality wireless data.
@@
Data augmentation is an effective technique to solve the problem of limited labeled datasets in model training[4]. The core idea of data augmentation is to synthetically increase the size of training datasets by modifying existing data, thus enhancing the robustness and generalization of learning algorithms[4]. Data augmentation techniques have been extensively applied in fundamental domains. For example, in Computer Vision (CV), data augmentation methods apply geometric or color transformations, such as cropping, flipping, and color channel changes[5]. However, traditional data augmentation methods may not be applicable to wireless data. Specifically, traditional data augmentation methods are carefully designed for certain domains and only perform simple transformations on the original datasets, rather than enriching data features from existing data[4]. Most importantly, traditional data augmentation methods lack consideration of the inherent characteristics and structures of wireless data[5], which cannot guarantee the quality and diversity of augmented wireless data. Hence, novel techniques for effective wireless data augmentation are urgently needed.
@@
Generative Artificial Intelligence (GenAI) is obtaining the full spotlight after the release of Large Language Model (LLM)-based chatbots by tech giants such as OpenAI’s ChatGPT, Google’s Bard, and Microsoft’s Bing Chat[6]. Unlike discriminative AI models that focus primarily on explicitly learning decision boundaries between classes, GenAI models excel at learning the underlying distribution, patterns, and features of input data, thus generating new data instances that resemble the original dataset[7]. Due to its transformative power in data generation, GenAI has recently gained significant attention in the realm of wireless networks[6], where real-world wireless data is often scarce, incomplete, costly to acquire, and difficult to model or comprehend[2], enabling the emergence of GenAI-driven data augmentation. Specifically, GenAI models can generate high-quality and diverse wireless data as a supplement, and the synthetic data can be combined with real data to augment the training dataset[6], which solves the wireless data acquisition challenge and can effectively improve the performance of DL models in wireless communication and networks.
@@
There are currently some preliminary studies on using GenAI models for data augmentation in wireless networks[2, 4, 8, 3]. For instance, the authors in [2] utilized conditional latent diffusion models to generate high-diversity and high-quality Radio Frequency (RF) data at low costs. In [3], the authors leveraged a Denoising Diffusion Probabilistic Model (DDPM) to generate channel data in multiple speed scenarios, where the DDPM can capture the underlying distribution of wireless channel data with limited data volume. Finally, a CsiNet model was used to validate the effectiveness of the proposed approach. However, the above works do not provide a general tutorial on how to implement data augmentation using GenAI technology in wireless networks. To this end, this paper focuses on answering two questions, i.e., “Why GenAI is conducive to data augmentation in wireless networks” and “How to use GenAI techniques to achieve wireless data augmentation.” To the best of our knowledge, this is the first work that systemically explores the potential and effectiveness of generative data augmentation in wireless networks. Our main contributions are summarized as follows:
@@
  • We review typical GenAI models and their applications in data augmentation, and summarize the benefits of generative data augmentation. We then explore the effectiveness of GenAI-driven data augmentation for wireless applications from the physical, network, and application layers, which presents a specific GenAI-driven data augmentation architecture for each application.
  • We propose a general Generative Diffusion Model (GDM)-based data augmentation framework for Wi-Fi gesture recognition, where we use transformer-based diffusion models to generate high-quality and diverse Channel State Information (CSI) data. We finally conduct a case study based on a real dataset, where we develop Residual neural Network (ResNet) models to evaluate the performance gains achieved by augmented CSI data. Simulation results demonstrate the effectiveness of the proposed framework.
@@ II-A1 Enhancing insufficient training datasets
Since wireless data, such as RF data, are sensitive and time-dependent to the open-space propagation environment[2], obtaining large amounts of high-quality training datasets for DL models in wireless communications is a costly and challenging endeavor. Fortunately, using effective data augmentation techniques can artificially increase the dataset size by generating variations of existing data, and subsequently mixing the existing training data with the augmented data into DL models, thus improving the performance and robustness of learning algorithms[4].
@@ II-A2 Boosting performance in security applications
In wireless security applications, anomaly detection is often affected by imbalanced datasets, where normal behavior far outweighs malicious events[9]. Especially, in an operational mechanical system that works under normal conditions most of the time, the collected sensor data representing positive training samples is sufficient[9], while machine operation failures are rare, and the corresponding collected sample is limited compared to positive samples. By using effective data augmentation techniques, training datasets with synthetic fault samples can improve the ability of DL models to detect rare events.
In the realm of machine learning and AI, several techniques are closely related to data augmentation but have distinct meanings, i.e., data synthesis and data enhancement. Specifically, data synthesis focuses on generating entirely new data instances that are not directly derived from existing data but created using GenAI models, while data enhancement is a complementary approach that aims to improve the quality of existing data, including augmented or synthesized data so that the data can be used more effective for training tasks.
@@ II-B Limitations of Non-AI Data Augmentation Techniques in Wireless Networks
Traditional (i.e., non-AI) data augmentation techniques resolve around simple transformations, such as cropping, rotations, and transformation[5]. Although traditional data augmentation techniques have been shown to significantly improve model performance, as wireless datasets grow in complexity, a series of intricate challenges are emerging:
  • Limited data diversity and balance: One of the challenges of data augmentation is to create enough diversity and balance in the data[3], especially for imbalanced datasets and Non-IID [^1] Non-IID means that data is not identically distributed. data. Data diversity refers to the variety and richness of data features, while data balance refers to the distribution and proportion of data labels. Classical data augmentation methods, such as rotation, flipping, and scaling, can generate only limited variations in the original data, which may not be sufficient to capture the variability and complexity of real-world data[5].
  • Domain-specific augmentation: Different domains have their own unique needs when it comes to data augmentation[4]. For instance, in medical imaging, the augmentation of medical images must be done cautiously to ensure the integrity of the data and avoid the introduction of artifacts that could compromise diagnosis. Thus, traditional augmentation techniques are often designed for specific types of data and are not universally applicable.
  • Loss of semantic information: Certain traditional data augmentation techniques may distort raw data in a way that alters semantic meanings[5]. For example, flipping an image horizontally may change the interpretation of content. In addition, the biases inherent in the original data are still present in the augmented data, and may even be reinforced.
@@ III-A Generative AI
  • Variational Autoencoders (VAEs): VAEs work by encoding the input data into a latent space and then decoding it back to the original space[7]. By sampling from decoding and the latent space, VAEs can generate new data that are closely consistent with the original distribution. In [10], VAEs are applied to augment intrusion detection data to enhance the performance of network security protection.
  • Generative Adversarial Networks (GANs): GANs consist of generator and discriminator networks that are trained simultaneously. The generator generates new data, while the discriminator tries to distinguish synthetic data from real data[7]. GANs are capable of dealing with the limitations of data expansion. For example, in [4], GANs are adopted for wireless channel data augmentation by learning the distribution of the original channel data.
  • Generative Diffusion Models: GDMs consist of the sampling procedure, the forward diffusion process, and the denoising process. The goal of GDMs is to learn to reverse the diffusion process to generate new data samples from the Gaussian noise[7]. GDMs are effective in data augmentation with the advantages of training straightforwardly and avoiding typical problems like model collapse[8]. For instance, in [8], GDMs are integrated with an attention module, which can generate high-quality and time-series RF data.
@@ III-B Benefits of Generative Data Augmentation
  • Privacy preservation: GenAI models can enable data augmentation without breaching privacy[6]. For example, GenAI models can generate synthetic datasets that resemble the original dataset without directly using sensitive data. Moreover, GenAI models can enhance federated learning systems by compensating for the missing portion of local data, where models are trained collaboratively without sharing raw data.
Figure 1: Overview of GenAI-driven data augmentation for wireless applications. We show how GenAI techniques can improve data augmentation to boost the performance of wireless applications at the physical, network, and application layers. In particular, we present a GenAI-driven data augmentation architecture for each application.
IV Generative AI-driven Data Augmentation for Wireless Applications
In this section, we explore how GenAI techniques play a role in data augmentation in wireless networks from the physical, network, and application layers. The overview of GenAI-driven data augmentation for wireless applications is shown in Fig. 1.
@@ IV-A Physical Layer
IV-A1 GenAI-driven data augmentation for wireless channel modeling
Wireless channel modeling involves simulating the propagation of electromagnetic waves between transmitters and receivers. However, collecting real-world channel data in different environments and under diverse conditions is often labor-intensive and has high overhead [4]. Fortunately, GenAI can generate a variety of synthetic channel responses, allowing for the creation of diverse training datasets that enhance the performance of DL-based wireless communication algorithms. For example, in [4], the authors developed a GAN-based network for channel data augmentation in an intelligent industrial wireless communication environment. Then, GAN-generated data is leveraged to train the DL-based channel state information feedback algorithm for performance enhancement. Simulation results demonstrate that for the accuracy of CSI feedback algorithms, the proposed GAN-based approach can obtain at most $3$ dB performance improvement compared with traditional data augmentation approaches.
IV-A2 GenAI-driven data augmentation for radio-frequency sensing
RF sensing is a field assisting vision technology that uses radio signals to capture surrounding information, enabling intelligent applications such as human activity recognition[2]. Due to the dynamic nature of the RF spectrum and the high cost of collecting RF datasets[2], GenAI models, especially GDMs, can be used to synthesize realistic RF data to augment existing datasets. For example, in [8], the authors proposed a time-frequency diffusion theory to instruct diffusion models to effectively generate time-series and high-quality RF data. The authors redesigned the denoising process by integrating attention-based diffusion blocks to extract RF features. Numerical results demonstrate the generation quality of the proposed diffusion model outperforms DDPMs by up to $42.4\%$.
@@ IV-B Network Layer
IV-B1 GenAI-driven data augmentation for network traffic management
@@
IV-B2 GenAI-driven data augmentation for network anomaly detection
Network anomaly detection is widely recognized as a crucial component of wireless network security. Traditional anomaly detection systems heavily rely on real-world data to learn the patterns of normal and abnormal behavior. However, gathering enough data to represent all possible exception types can be challenging. Generative data augmentation provides a feasible solution for enhancing network anomaly detection. For example, in [12], the authors proposed a GAN-based method called TAnoGan for unsupervised anomaly detection when time series data points are small. Numerical results demonstrate that the proposed TAnoGan ranks higher than baseline models in $F_{1}$ score, which is the most important indicator in anomaly detection [12].
Figure 2: A GDM-based data augmentation framework for Wi-Fi gesture recognition. The framework consists of three parts: Part 1 is the pre-processing raw CSI data; Part 2 is the use of transformer-based diffusion models to synthesize high-quality CSI data; Part 3 is the evaluation of augmented data.
@@ IV-C Application Layer
IV-C1 GenAI-driven data augmentation for object detection
In wireless networks, object detection tasks mainly involve identifying and tracking the location of users or devices. However, object detection requires not only target labels in each image but also precise bounding boxes[13]. This extra work makes the management of these datasets for training object detection models more laborious than image classification[13]. An alternative approach to effectively manage datasets for object detection is generative data augmentation. For example, the authors in [13] proposed a data augmentation pipeline based on contrastive language-image pertaining and controllable diffusion models, which can generate images with high-quality bounding box annotations.
IV-C2 GenAI-driven data augmentation for digital twins

### v2

@@ Abstract
Data augmentation as a technique can mitigate data scarcity in machine learning. However, owing to fundamental differences in wireless data structures, traditional data augmentation techniques may not be suitable for wireless data. Fortunately, Generative Artificial Intelligence (GenAI) can be an effective solution to wireless data augmentation due to its excellent data generation capability. This article systematically explores the potential and effectiveness of generative data augmentation in wireless networks. We first briefly review data augmentation techniques, discuss their limitations in wireless networks, and introduce generative data augmentation, including reviewing GenAI models and their applications in data augmentation. We then explore the application prospects of generative data augmentation in wireless networks from the physical, network, and application layers, providing a generative data augmentation architecture for each application. Subsequently, we propose a general generative data augmentation framework for Wi-Fi gesture recognition. Specifically, we leverage transformer-based diffusion models to generate high-quality channel state information data. To evaluate the effectiveness of the proposed framework, we conduct a case study using the Widar 3.0 dataset, which employs a residual network model for Wi-Fi gesture recognition. Simulation results demonstrate that the proposed framework can enhance the performance of Wi-Fi gesture recognition. Finally, we discuss research directions for generative data augmentation.
@@
Wireless networks, data augmentation, GenAI, Wi-Fi gesture recognition, transformer-based diffusion models.
@@ I Introduction
As an indispensable and fundamental technology, wireless networks enable users to access the Internet, connect devices, and communicate without physical constraints. The growing complexity and diversity of wireless networks have facilitated the development of Deep Learning (DL)-based wireless communications and networking [1]. Since wireless networks generate large amounts of data from user activities, channel states, and network conditions, DL can efficiently distill high-level features and information from this data, which have complex structures and inner correlations [1], and enable intelligent applications such as Digital Twins (DTs), autonomous driving, and Metaverse. Nevertheless, the availability of vast amounts of high-quality wireless data is a major determinant of the effectiveness of DL models [2]. Specifically,
  • Dynamic wireless channels: Due to multi-path fading, shadowing, and interference, the environment in wireless channels is highly variable [3], which makes it difficult to collect high-quality wireless data.
@@
Data augmentation is an effective technique to solve the problem of limited labeled datasets in model training [4]. The core idea of data augmentation is to synthetically increase the size of training datasets by modifying existing data, thus enhancing the robustness and generalization of learning algorithms [4]. Data augmentation techniques have been extensively applied in fundamental domains. For example, in Computer Vision (CV), data augmentation methods apply geometric or color transformations, such as cropping, flipping, and color channel changes [5]. However, traditional data augmentation methods may not be applicable to wireless data. Specifically, traditional data augmentation methods are carefully designed for certain domains and only perform simple transformations on the original datasets, rather than enriching data features from the existing data [4]. Most importantly, traditional data augmentation methods lack consideration of the inherent characteristics and structures of wireless data [5], which cannot guarantee the quality and diversity of augmented wireless data. Hence, novel techniques for effective wireless data augmentation are urgently needed.
@@
Generative Artificial Intelligence (GenAI) is obtaining the full spotlight after the release of Large Language Model (LLM)-based chatbots by tech giants such as OpenAI’s ChatGPT, Google’s Bard, and Microsoft’s Bing Chat [6]. Unlike discriminative AI models that focus primarily on explicitly learning decision boundaries between classes, GenAI models excel at learning the underlying distribution, patterns, and features of input data, thus generating new data instances that resemble the original dataset [7]. Due to its transformative power in data generation, GenAI has recently gained significant attention in the realm of wireless networks [6], where real-world wireless data is often scarce, incomplete, costly to acquire, and difficult to model or comprehend [2], enabling the emergence of generative data augmentation. Specifically, GenAI models can generate high-quality and diverse wireless data as a supplement, and the synthetic data can be combined with real data to augment the training dataset [6], which solves the wireless data acquisition challenge and can effectively improve the performance of DL models in wireless communication and networks.
@@
There are currently some preliminary studies on using GenAI models for data augmentation in wireless networks [2, 4, 8, 3]. For instance, the authors in [2] utilized conditional latent diffusion models to generate high-diversity and high-quality Radio Frequency (RF) data at low costs. In [3], the authors leveraged a Denoising Diffusion Probabilistic Model (DDPM) to generate channel data in multiple speed scenarios, where the DDPM can capture the underlying distribution of wireless channel data with limited data volume. Finally, a CsiNet model was used to validate the effectiveness of the proposed approach. However, the above works do not provide a general tutorial on how to implement generative data augmentation in wireless networks. To this end, this paper focuses on answering two questions, i.e., “Why GenAI is conducive to data augmentation in wireless networks?” and “How to use GenAI techniques to achieve wireless data augmentation?To the best of our knowledge, this is the first work that systematically and comprehensively explores the applications of generative data augmentation in wireless networks. Our main contributions are summarized as follows:
@@
  • We first review typical GenAI models and their applications in data augmentation, and summarize the benefits of generative data augmentation. We then systematically and comprehensively explore the effectiveness of generative data augmentation for wireless applications from the physical, network, and application layers, which presents a specific generative data augmentation architecture for each wireless application.
  • We propose a general generative data augmentation framework for Wi-Fi gesture recognition. This framework leverages transformer-based diffusion models, which represent the State-of-the-Art (SOTA) model in RF data augmentation, to generate high-quality and diverse Channel State Information (CSI) data. To evaluate the effectiveness of the proposed framework, we conduct a case study using the Widar 3.0 dataset, which employs a Residual Network (ResNet) model for Wi-Fi gesture recognition. Extensive simulation results demonstrate the effectiveness of the proposed framework.
@@ II-A1 Enhancing insufficient training datasets
Since wireless data, such as RF data, is sensitive and time-dependent to the open-space propagation environment [2], obtaining large amounts of high-quality training datasets for DL models in wireless communications is a costly and challenging endeavor. Fortunately, using effective data augmentation techniques can artificially increase the dataset size by generating variations of existing data, and subsequently mixing the existing training data with the augmented data into DL models, thus improving the performance and robustness of learning algorithms [4].
@@ II-A2 Boosting performance in security applications
In wireless security applications, anomaly detection is often affected by imbalanced datasets, where normal behavior far outweighs malicious events [9]. For instance, in operational mechanical systems that work under normal conditions most of the time, the collected sensor data representing positive training samples is sufficient [9], while machine operation failures are rare, and the corresponding collected sample is limited compared to positive samples. By using effective data augmentation techniques, training datasets with synthetic fault samples can improve the ability of DL models to detect rare events.
In the realm of machine learning and AI, several techniques are closely related to data augmentation but have distinct meanings, i.e., data synthesis and data enhancement. Specifically, data synthesis focuses on generating entirely new data instances that are not directly derived from existing data but created using GenAI models, while data enhancement is a complementary approach that aims to improve the quality of existing data, including augmented or synthesized data, so that the data can be used more effectively for training tasks.
@@ II-B Limitations of Non-AI Data Augmentation Techniques in Wireless Networks
Traditional (i.e., non-AI) data augmentation techniques primarily rely on simple transformations such as cropping, rotations, and transformation [5]. Although traditional data augmentation techniques have been shown to significantly improve model performance, as wireless datasets grow in complexity, a series of intricate challenges are emerging.
  • Limited data diversity and balance: One of the challenges of data augmentation is to create enough diversity and balance in the data [3], especially for imbalanced datasets and Non-IID [^1] Non-IID means that data is not identically distributed. data. Data diversity refers to the variety and richness of data features, while data balance refers to the distribution and ratio of data labels. Classical data augmentation methods, such as rotation, flipping, and scaling, can generate only limited variations in the original data, which may not be sufficient to capture the variability and complexity of real-world data [5].
  • Domain-specific augmentation: Different domains have their own unique needs when it comes to data augmentation [4]. For instance, in medical imaging, the augmentation of medical images must be done cautiously to ensure the integrity of the data and avoid the introduction of artifacts that could compromise diagnosis. Thus, traditional data augmentation techniques are often designed for specific types of data and are not universally applicable.
  • Loss of semantic information: Traditional data augmentation techniques may distort raw data in ways that alter semantic meanings [5]. For example, flipping images horizontally may change the content interpretation. Moreover, biases inherent in the original data are still present in the augmented data and may even be reinforced.
@@ III-A Generative AI
  • Variational Autoencoders (VAEs): VAEs work by encoding the input data into a latent space and then decoding it back to the original space [7]. By sampling from decoding and the latent space, VAEs can generate new data that is closely consistent with the original distribution. In [10], VAEs are applied to augment intrusion detection data to enhance the performance of network security protection.
  • Generative Adversarial Networks (GANs): GANs consist of a generator and a discriminator, which are trained simultaneously in a minimax game framework. The generator generates new data, while the discriminator tries to distinguish synthetic data from real data [7]. GANs can deal with the limitations of data expansion. In [4], GANs are adopted for wireless channel data augmentation by learning the distribution of the original channel data.
  • Generative Diffusion Models (GDMs): GDMs consist of three key components: a sampling procedure, a forward diffusion process, and a denoising process. The goal of GDMs is to learn to reverse the diffusion process to generate new data samples from the Gaussian noise [7]. GDMs are effective in data augmentation with the advantages of training straightforwardly and avoiding typical problems like model collapse [8]. For instance, in [8], GDMs are integrated with an attention module, which can generate high-quality and time-series RF data.
@@ III-B Benefits of Generative Data Augmentation
  • Privacy preservation: GenAI models can enable data augmentation without breaching privacy [6]. For example, GenAI models can generate synthetic datasets that resemble the original dataset without directly using sensitive data. Moreover, GenAI models can enhance federated learning systems by compensating for the missing portion of local data, where models are trained collaboratively without sharing raw data, avoiding privacy leakage issues.
Figure 1: Overview of generative data augmentation for wireless applications. We show how GenAI techniques can improve data augmentation to enhance the performance of wireless applications at the physical, network, and application layers. In particular, we present a generative data augmentation architecture for each application.
IV Generative Data Augmentation for Wireless Applications
In this section, we explore how GenAI techniques play a role in data augmentation in wireless networks from the physical, network, and application layers, as shown in Fig. 1.
@@ IV-A Physical Layer
IV-A1 Wireless channel modeling
Wireless channel modeling involves simulating the propagation of electromagnetic waves between transmitters and receivers. However, collecting real-world channel data in different environments and under diverse conditions is often labor-intensive and has high overhead [4]. Fortunately, GenAI can generate a variety of synthetic channel responses, allowing for the creation of diverse training datasets that enhance the performance of DL-based wireless communication algorithms. For example, in [4], the authors developed a GAN-based network for channel data augmentation in an intelligent industrial wireless communication environment. Then, GAN-generated data is leveraged to train the DL-based channel state information feedback algorithm for performance enhancement. Simulation results demonstrate that for the accuracy of CSI feedback algorithms, the proposed GAN-based approach can obtain at most $3\>\rm{dB}$ performance improvement compared with traditional data augmentation approaches.
IV-A2 Radio-frequency sensing
RF sensing is a field assisting vision technology that uses radio signals to capture surrounding information, enabling intelligent applications such as human activity recognition [2]. Due to the dynamic nature of the RF spectrum and the high cost of collecting RF datasets [2], GenAI models, especially GANs and GDMs, can be used to synthesize realistic RF data to augment existing datasets. For example, in [8], the authors proposed a time-frequency diffusion theory to instruct diffusion models to effectively generate time-series and high-quality RF data. The authors redesigned the denoising process by integrating attention-based diffusion blocks to extract RF features. Simulation results demonstrate that the generation quality of the proposed diffusion model outperforms DDPMs by up to $42.4\%$.
@@ IV-B Network Layer
IV-B1 Network traffic management
@@
IV-B2 Network anomaly detection
Network anomaly detection is a crucial component of wireless network security. Traditional anomaly detection systems rely heavily on real-world data to learn the patterns of normal and abnormal behavior. However, gathering enough data to represent all possible exception types can be challenging. Generative data augmentation provides a feasible solution for enhancing network anomaly detection. For example, in [12], the authors proposed a GAN-based method for unsupervised anomaly detection when time series data points are small. Simulation results demonstrate that the proposed method ranks higher than baseline models in $F_{1}$ score, which is the most important indicator in anomaly detection [12].
Figure 2: A generative data augmentation framework for Wi-Fi gesture recognition. The framework consists of three parts: Part 1 is the pre-processing of raw CSI data; Part 2 is the use of transformer-based diffusion models to synthesize high-quality CSI data; Part 3 is the evaluation of augmented data. Note that the proposed framework can also be applied to other RF data, and the corresponding code is available in https://github.com/JinboWen00/GDA-for-Wi-Fi-Gesture-Recognization.
@@ IV-C Application Layer
IV-C1 Object detection
In wireless networks, object detection tasks mainly involve identifying and tracking the location of users or devices. However, object detection requires not only target labels in each image but also precise bounding boxes [13]. This extra work makes the management of these datasets for training object detection models more laborious than image classification [13]. An alternative approach to effectively managing datasets for object detection is generative data augmentation. For example, the authors in [13] proposed a data augmentation pipeline based on contrastive language-image pertaining and controllable diffusion models, which can generate images with high-quality bounding box annotations.
IV-C2 Digital twins